### v2

# Nuclear Spin-Mediated Relaxation Mechanisms of the V Center in hBN

Abstract
The negatively charged boron vacancy (V) defect in hexagonal boron nitride has recently emerged as a promising spin qubit for sensing due to its high-temperature spin control and versatile integration into van der Waals structures. While extensive experiments have explored their coherence properties, much less is known about the spin relaxation time () and its control parameter dependence. In this work, we develop a parameter-free spin dynamics model based on the cluster expansion technique to investigate relaxation mechanisms at low temperature. Our results reveal that the V center constitutes a strongly coupled electron spin-nuclear spin core, which necessitates the inclusion of the coherent dynamics and derived memory effects of the three nearest-neighbor nitrogen nuclear spins. Using this framework, this work closely reproduces the experimentally observed time at G and further predicts the dependence on external magnetic field in the G interval, when the spin relaxation is predominantly driven by electron–nuclear and nuclear–nuclear flip–flop processes mediated by hyperfine and dipolar interactions. This study establishes a reliable and scalable approach for describing relaxation in V centers and offers microscopic insights to support future developments in nuclear-spin-based quantum technologies.
Department of Computer Engineering, TUM School of Computation, Information and Technology, Technical University of Munich, 80333 Munich, Germany
Munich Center for Quantum Science and Technology (MCQST), 80799 Munich, Germany
Abbe Center of Photonics, Institute of Applied Physics, Friedrich Schiller University Jena, 07745 Jena, Germany
Department of Physics of Complex Systems, Eötvös Loránd University, Egyetem tér 1-3, H-1053 Budapest, Hungary
MTA–ELTE Lendület ”Momentum” NewQubit Research Group, Pázmány Péter, Sétány 1/A, 1117 Budapest, Hungary
email: tobias.vogl@tum.de, ivady.viktor@ttk.elte.hu
November 21, 2025
1 Introduction
In solid-state quantum systems, spin qubits have emerged as a crucial resource for advancing quantum technologies, including quantum sensing[10.1038/s41563-024-01887-z, 10.1038/s41467-021-24725-1], quantum memory[10.1103/PhysRevX.6.021040, 10.1038/nphys2026, 10.1126/science.1220513, 10.1063/5.0188597, 10.1002/adom.202302760], quantum computing[10.1063/5.0007444], and quantum repeaters[10.1038/s41534-022-00637-w]. Among various defect-based platforms, the nitrogen-vacancy (NV) center in diamond has demonstrated remarkable potential, owing to its long spin coherence times and the ability to initialize, coherently manipulate, and optically read out individual electron spins at room temperature[10.1016/j.physrep.2013.02.001, 10.1126/science.1139831]. Motivated by these achievements, considerable effort has recently focused on identifying and characterizing other optically addressable spin defects in wide-bandgap materials[10.1038/s41563-024-01887-z, 10.1021/acsnano.3c08940, 10.1038/s41467-024-51129-8, 10.1021/acs.jpclett.3c01475]. In particular, point defects in hexagonal boron nitride have attracted growing interest due to their intrinsically high photon out-coupling efficiency[10.1038/nnano.2015.242], excellent compatibility with photonic components[10.1088/1361-6463/aa7839, 10.1002/adom.202002218], long-term photostability[10.1021/acsphotonics.8b00127, 10.1016/j.physe.2020.114251], bright and spectrally pure single-photon emission[10.1021/acsnano.3c08940], as well as various defect choices[10.3390/nano12142427, 10.1021/acs.jpcc.4c03404].
Among these, the negatively charged boron vacancy defect has emerged as a candidate for enabling coherent control of an electron spin and nearby nuclear spins in hBN. The V center consists of a missing boron atom coordinated by three nitrogen atoms in a D3h symmetry, forming an electronic spin system with a zero-field splitting (ZFS) of approximately 3.5 GHz [10.1038/s41467-023-44494-3, 10.1038/s41563-020-0619-6, 10.1038/s41524-020-0305-x]. Recent experimental studies have characterized its spin and optical properties [10.1038/s41467-023-44494-3, 10.1038/s41563-020-0619-6, 10.1038/s41524-020-0305-x, 10.1126/sciadv.abf3630, 10.1021/acs.nanolett.1c02495, 10.1038/s41563-022-01329-8]. At room temperature, longitudinal relaxation times have been reported in the range of 15–18 s, largely independent of applied magnetic field and host isotope composition [10.1038/s41467-023-44494-3, 10.1126/sciadv.abf3630, 10.1021/acs.nanolett.1c02495]. At cryogenic temperatures, extends to 1 ms up to 12.5 ms, depending on the isotope concentration and the magnetic field strength[10.1038/s41467-023-44494-3, 10.1126/sciadv.abf3630]. Recent experiments have demonstrated that the V center remains stable in ultrathin, few-layer hBN flakes, exhibiting robust optically detected magnetic resonance signals and spin polarization, while spin-lattice relaxation appears to depend on the dimensionality of the sample.[durand_optically_2023] However, the comprehensive analysis about dependence of on the external magnetic field remains largely unexplored.
In terms of spin coherence, spin echo times () are strongly influenced by the nuclear spin environment, yielding values of 82.1 ns for natural isotopic abundance hBN[10.1038/s41467-022-33399-2], 46 ns for h11BN, and 62 ns for h10BN[10.1038/s41467-022-31743-0]. While most early investigations focused on 14N-containing systems, recent work has identified hB15N as a particularly favorable isotopic configuration, offering improved and times due to reduced hyperfine interactions[10.1038/s41467-023-44494-3, 10.1002/adfm.202511300, 10.1038/s41699-022-00336-2].
Despite significant experimental and theoretical progress in understanding decoherence in V center[10.1038/s41467-023-44494-3, 10.1038/s41467-022-31743-0, rizzato_extending_2023, ramsay_coherence_2023, 10.1002/adfm.202511300, 10.1038/s41699-022-00336-2], comprehensive theoretical modeling of their spin relaxation dynamics remains limited. In particular, while several computational studies have focused on investigating time in these systems[10.1002/adfm.202511300, 10.1038/s41699-022-00336-2, 10.1103/PhysRevB.106.104108], systematic modeling of the time has been comparatively scarce. A thorough understanding of the mechanisms governing is essential, as it ultimately sets an upper bound for the achievable spin coherence time through the fundamental relation [10.1088/0034-4885/80/1/016001]. Typically, the relaxation time comprises multiple contributions arising from distinct microscopic processes, including spin–spin and spin–lattice interactions[10.1126/sciadv.abf3630, 10.1103/PhysRevB.97.094304, 10.1038/s41524-021-00673-8]. At room temperature, spin relaxation is predominantly governed by spin-lattice interactions[10.1038/s41467-023-44494-3, 10.1103/PhysRevLett.108.197601]. This can be theoretically captured by ab initio spin dynamics simulations[10.1038/s41524-023-01082-9, 10.1103/PhysRevB.98.214442]. However, at cryogenic temperatures, where phonon populations are suppressed, spin relaxation is largely dictated by spin flip-flop processes through hyperfine and dipolar interactions with the surrounding nuclear spin bath. Capturing dipolar spin relaxation is particularly challenging, as the approximations used for computing decoherence in cluster correlation expansion methods are not applicable, and many-spin entanglements may need to be taken into account, especially in the dense nuclear spin environment in hBN[10.1103/PhysRevB.101.155203]. To date, no comprehensive model has been constructed to fully capture these low-temperature spin relaxation mechanisms in V centers.
In this work, we therefore develop a spin dynamics model to investigate the relaxation mechanism of V centers in bulk hBN at low temperature and moderate magnetic fields. To accurately account for the interplay of hyperfine interactions, dipolar couplings, and electron-nuclear entanglement effects, we employ a cluster expansion technique combined with an extended Lindbladian formalism[10.1103/PhysRevB.101.155203, xu_ab_2024]. This approach enables us to approximate the time evolution of the diagonal elements of the reduced density matrix of the central spin system without simplifying the underlying spin interactions, while remaining computationally tractable. Using this framework, we first carry out various models and then identify the most effective model capable of accurately capturing spin decoherence. We show that the electron spin-three first-neighbor 15N nuclear spin coupled system must be treated as a single coherent unit, an extended core, in order to reproduce the experimentally observed decay characteristics. We then apply this optimized model to elucidate the relaxation mechanisms responsible for the experimentally observed at G[10.1038/s41467-023-44494-3]. Finally, we extend our analysis to predict the dependence on the external magnetic field. These results offer valuable microscopic insight into the spin relaxation dynamics of V centers and establish a theoretical foundation for future developments in coherence control, defect engineering, and device optimization for quantum information applications.
2 Results
The analysis of relaxation time is structured into three different subsections. First, we investigate the variations in the values derived from four computational models, which include memory effects of the bath on different levels. The relevance of this part is to gain deeper insight into the decay mechanism of the V center and to identify a suitable, feasible model for computing . Next, we assess predictions of the optimal model against the experimental measurements. Finally, we examine the dependence of time on the external strength.
To investigate the spin relaxation mechanism of the V defect in h11B15N bulk under low temperature, we focus on spin relaxation induced by interactions between the electron of the spin of the V and the nuclear spin bath, as depicted in Fig 2(a). These interactions can be described by the following spin Hamiltonian
where accounts for the Hamiltonian for the central spin; represents the Hamiltonian for the bath spins in the environment; and represents the Hamiltonian for the interactions between the electron spin and the bath spins through the hyperfine (HF) interaction. The electron spin Hamiltonian can be expressed as
where the first term on the right-hand side of the equation describes the electron Zeeman interaction, while the second and third terms describe the zero-field splitting interaction. Here, is Bohr magneton, is an external magnetic field in the direction perpendicular to the hBN, is a factor of an electron, for an electronic spin-1, is the zero-field splitting equal to 3.479 GHz[10.1038/s41524-020-0305-x] at room temperature, and is the transverse component of the ZFS, which is set to zero in this work. We note that the unit of frequency (Hz), and not the angular frequency (rad-Hz), is used throughout this paper.
The electron spin Zeeman and zero-field splitting terms of are the leading terms in Eq. 1, which determine the main characteristics of the spectrum of the spin systems and have a major effect on the relaxation processes. At zero magnetic field, the eigenstates corresponding to and are split in the ground state due to the zero-field splitting. When an external magnetic field is applied, these eigenstates undergo additional splitting through the Zeeman interaction, resulting in the separation of the states, as shown in Fig. 1.
The Hamiltonian of the bath contains the following interactions
where is nuclear spin operator vector, is a gyromagnetic ratio of the nuclear spins, and is the number of nuclear spins considered. Here, the first term accounts for nuclear Zeeman interaction, whereas the second term accounts for the dipole-dipole nuclear interaction, which is responsible for the nuclear-nuclear spin flip-flop process.
Finally, the interaction between the central spin and the spin bath can be expressed as
where is the hyperfine-interaction tensor, calculated based on the method in Ref. [10.1038/s42005-024-01668-9] to tackle the finite-size effect. It is important to note that each branch in Fig. 1 is further divided into four sublevels due to hyperfine interactions with the three nuclear spins. These hyperfine splittings are approximately 66 MHz for 15N. At a magnetic field of approximately 1250 G, the eigenenergies of the and states approach near-degeneracy, see inset of Fig. 1. In this regime, strong mixing occurs between the electronic and nuclear spin states, enabling efficient spin flip-flop processes. This enhanced flip-flop interaction leads to pronounced correlations between spins, consequently modifying the T1 relaxation behavior, as will be discussed later.
As shown in Eq. (4), the , , , and hyperfine components govern the electron–nuclear spin flip-flop processes. To illustrate the strength of these interactions, their values are plotted in Fig. 2(b), highlighting the strongly coupled nature of the V defect system. Fig. 2(b) unravels that rotational invariant perpendicular hyperfine components of the first three nearest nitrogen nuclei are strong due to the Fermi contact contribution. This suggests that relaxation via the electron–nuclear spin flip-flop process in this channel is indispensable and plays a central role in the overall spin dynamics.
To simulate spin relaxation process, our models take into account all major parts of the spin Hamiltonian in Eq. (1) and truncate only the nuclear spin-nuclear spin dipolar couplings to a certain degree depending on the clustering used. The dimension of the given total Hamiltonian depends on the dimension of the Hilbert space of electron and nuclear spins, and the computational demand increases exponentially with the number of nuclear spins, limiting the capability of the exact solution. To resolve this issue, the methodology developed in Ref. [10.1103/PhysRevB.101.155203] is applied by dividing the entire system into smaller clusters of spins utilizing the central spin approximation and cluster expansion to reduce the Hilbert space considered while still maintaining all the aforementioned interactions.
In short, the Hamiltonian of the cluster of the spin system can be written as
We simulate the dynamics of the system through the density matrix () following the below Master equation with an extended Lindbladian developed by Ref. [10.1103/PhysRevB.101.155203]
We note that represents an extended Lindbladian term introduced into the Master equation to account for spin flip-flop processes induced by environmental spins not included in cluster (diagonal elements in the density matrix), see Ref. [10.1103/PhysRevB.101.155203] for more details. Additionally, it is used to ensure the synchronization of key properties of different instances of the central spin across different clusters, such as expectation values. This correction is essential because the unsynchronization of a central spin in all clusters can lead to an artificial thermal state due to the accumulation of its magnetization. The extended term can be formulated as follows
Here, determines spin flip-flop rates of the central spin induced by environmental spins not incorporated in cluster and is the jump operator describing the spin flip and flop transitions of the central spin between and states. As such, Eq. (7) not only causes decoherence but also contributes to the decay of the central spin.
Even though the electron spin-nuclear interactions are not approximated, the Hilbert space is significantly reduced through clusterization. In this approach, the subsystems of the central spin-spin bath model consist of only a few spins at a time, which limits the entanglement and non-classical correlations in the system that can be captured. Consequently, the choice of clusters and the definition of subsystems directly determine the accuracy of the approximation. For strongly interacting systems such as the V defect in hBN, there is no established guideline for constructing such clusterizations. The cluster must be defined in a way that ensures the density matrix remains computationally tractable while accurately capturing the relevant correlation effects. Therefore, we investigate four different models with varying complexity and subsystem sizes to uncover relevant correlations and memory effects within the bath.
2.1 Theoretical Prediction
As detailed in the Methods section, the time is determined from the population of the electron spin until it relaxes to thermal equilibrium. Fig. 1 illustrates the energy-level splitting among the three relevant states and suggests that the fields in the vicinity of the ground-state level anticrossing (GSLAC) at around 1250 G yield particularly narrow splitting, as the GSLAC is the point at which two ground-state sublevels become nearly degenerate. This implies that spin relaxation near GSLAC is faster, which can minimize computational expense. We therefore adopt a fixed field at 800 G to evaluate the performance of various computational models for predicting . This value is close to, but outside, the GSLAC regime.
2.2 Model 1: Two-Spin Clusters
In this model, each cluster consists of the electron spin-1 of V and a single nitrogen nuclear spin, as illustrated in Fig. 3(a). This results in a Hilbert space of dimension , corresponding to six basis states. In total, 32 subsystems have been constructed in the cluster. The boron nuclear spins are excluded due to their weaker hyperfine interactions at this point, as shown in Fig. 2(b). Based on this cluster, the dipolar nuclear interaction between nuclear spins is neglected. In addition, due to the two-spin clusterization, only electron spin-nuclear spin correlation and coherence are accounted for by the approximation.
Fig. 4(a) depicts the evolution of the V electron spin initialized in the ms = 0 state. The result suggests that immediately after initialization, the electron spin polarization is rapidly transferred to the 15N nuclear spin in this model. At first, the relaxation appears exponential, but as the m population approaches the thermal equilibrium value of 1/3, as expected for the equal occupancy of the three spin sublevels (m), the relaxation pattern deviates from the exponential curve[10.1103/PhysRevB.65.205309, 10.1126/science.1155400] and exhibits a stretched exponential decay. This is confirmed by the exponent value of 0.37. Stretched exponents often arise when the relaxation is governed by multiple, independent decay mechanisms with different rates. While the distribution of the perpendicular hyperfine components, comprising strongly, moderately, and weakly coupled spins, see Fig. 2(b), could suggest such a multi-channel relaxation mechanism, experimental results do not support this model. As a consequence, we conclude that Model 1 fails to capture relevant correlation effects of the bath that may lead to different decay mechanisms.
2.3 Model 2 - Three/Four-Spin Clusters
In this model, clusters consisting of nitrogen and boron nuclei are constructed, as depicted in Fig. 3(b). Each nitrogen cluster includes an electron spin-1 coupled with three nitrogen nuclear spins, whereas each boron cluster consists of an electron spin-1 coupled with only two boron nuclear spins to compensate for the larger boron spin.
To reduce the computational complexity of the density matrix, we apply a reduced density matrix approach to the electron spin-1 system. Specifically, we restrict the spin subspace to m and ms = -1, while the ms = 1 state is truncated. This approximation retains the essential spin interactions close to GSLAC while significantly (by ) reducing the density matrix size.
For nitrogen clusters, the Hilbert space dimension is
while for boron clusters, the Hilbert space dimension is
Unlike Model 1, this model incorporates dipolar interactions within each cluster, which may additionally contribute to spin dynamics behavior. In addition, Model 2 accounts for up to 4 spin correlation effects, thus, it significantly improves on Model 1 and can describe more complex dynamics. Here, 47 nitrogen clusters and 48 boron clusters have been used, describing a 238-spin model in total. Note that the number of clusters has been tested as shown in Supplementary S2.
Fig. 4(b) illustrates that the electron‐spin polarization continues to be rapidly transferred to the surrounding nuclear spins until an equilibrium population. Note that the equilibrium population reaches a value of 0.5 because the reduced density matrix is restricted to the m and m sublevels. Furthermore, although the relaxation decay profile appears somewhat closer to exponential than that observed in Model 1, it nonetheless deviates from the experimental exponential behavior. These observations indicate that, despite the incorporation of nuclear dipole–dipole interactions and increasing the Hilbert space of the cluster systems, the present model does not fully account for all effects necessary to reproduce the anticipated decay dynamics. Thus, an even more comprehensive modeling approach is required.
2.4 Model 3 - Three/Four-Spin Clusters with Additional Dephasing
Before moving toward another clusterization, we investigate the effect of decoherence of the electron spin on time. The cluster Hamiltonian in this model remains the same as in Model 2, and the number of clusters is also identical. However, in this model, we incorporate the effects of spin dephasing by introducing a dephasing time () of 200 ns [10.1038/s41467-023-44494-3] into the off-diagonal terms of the Lindbladian in Eq. (7). Note that the variation of values has been tested in Supplementary S3.
Physically, characterizes the timescale over which phase coherence between quantum states is lost due to internal noise sources. By incorporating dephasing, modeled here as an exponential decay of the magnitude of the off-diagonal elements of the reduced density matrix of the electron spin, we suppress coherent oscillations arising from Larmor precession and hyperfine-driven spin flip-flops. This, in turn, enhances the overall relaxation dynamics and shortens the .
As shown in Fig. 4(c), the relaxation rate () increases by an order of magnitude as expected, however, the overall decay profile remains a stretched exponential and exhibits a lower decay exponent () compared to Model 2. These findings imply that the introduced dephasing term partially restores the behavior observed in Model 1, but it fails to accurately capture the correct relaxation dynamics.
2.5 Model 4 - Six-Spin Clusters
In this model, each cluster consists of an electron spin-1 and the three nearest-neighbor nitrogen nuclei, giving rise to an extended 4-spin central spin model, and two additional distant nitrogen nuclei, whose different in each cluster, as shown in Fig. 3(c). Consequently, the Hilbert space dimension is given by
These central and bath spin configurations enhance both the hyperfine coupling and the dipolar interaction terms in the Hamiltonian. This model accounts for all the correlation effects between the electron spin and the first neighbor nitrogen nuclear spins, as well as correlations and interactions with two additional nuclear spins.
To manage computational complexity, we apply the same reduced density matrix approach as in the previous two models. Furthermore, a total of 28 clusters are considered in this model, accounting for a total of 60 spins. It is important to note that in the model, we focus on nitrogen nuclear spins, and boron nuclei are neglected at first, as they exhibit a weak hyperfine interaction. Furthermore, previous studies have shown that heteronuclear spin flip-flop processes, e.g., 29Si-13C flip-flops in silicon carbide and 11B-15N flip-flops in hBN, are strongly suppressed at nonzero fields due to the differences in the gyromagnetic ratios of the nuclear spins.[10.1103/PhysRevB.90.241203] Consequently, in heterogeneous clusters consisting of the V electron spin, the three nitrogen nuclear spins in the central spin, and additional boron atoms, no direct coupling between the two types of nuclear spins is expected. We therefore assume that the nitrogen and boron nuclear baths decouple from each other, analogous to the case of decoherence.[10.1002/adfm.202511300] We note, however, that in contrast to SiC, the nuclear spins in hBN are much more strongly coupled to the electron spin, which may lead to electron-spin-mediated nuclear spin flip-flops[10.1002/adfm.202511300] and potentially to additional spin relaxation mechanisms. Therefore, the computed time obtained from Model 4 should be regarded as an upper bound. Note that the timestep convergence test and the effect of nuclear-spin synchronization have been validated in Supplementary S4.
Model 4 enlarges the central‐spin subsystem to include the V electron spin and its three nearest‐neighbor 15N nuclei, while the bath is augmented by two additional distant 15N nuclei. The initial electron–spin polarization in the ms = 0 sublevel is set to 70%, in agreement with experimental measurements[10.1002/qute.202300118, 10.1103/PhysRevB.110.014104]. Note that the previous models initialized the electron spin polarization to 100% in the ms = 0 state for the sake of identifying the proper model. Since Model 4 includes all relevant interactions, we initialize the electron spin according to the experimental value of 70%. Under these conditions, the expected exponential relaxation of electron spin polarization appears, see Fig. 4(d). The decay behavior is well described by an exponential function (decay exponent equal to 1). This demonstrates that the previously unaccounted correlation effects arise from insufficient entanglement and dipolar interactions within the extended central spin and also with the bath spins. That is, the spin dynamics are mediated by the spin flip-flop processes. Specifically, referring to Eqs. (3) and (4), nuclear–nuclear flip–flop transitions are driven by the dipolar interaction term whereas electron–nuclear flip–flop transitions originate from the non‐secular hyperfine term . As illustrated in Fig. 2(b), the V electron spin exhibits strong hyperfine coupling to its three nearest nitrogen nuclei, therefore, including these four spins in the central‐spin description is essential to accurately capture the unseparable dynamics of the electron spin and the first neighbor nucelar spins. Simultaneously, the addition of two distant nuclear spins enables an accurate description of nuclear–nuclear flip‐flops via the dipolar interaction and additional correlation effects. Model 4 can be considered as defining the boundary between the central spin and the spin bath at the first-neighbor shell of the V center. Consequently, the reduced bath includes only the moderately and weakly coupled nuclear spins. In this extended central spin–moderately coupled spin bath model, the spin bath can be considered “memoryless”, or at least its correlations have no significant influence on the dynamics of the extended central spin system, which, in turn, gives rise to an exponential decay characteristic of Markovian processes. As a result, this model encompasses all significant relaxation pathways accurately and will be employed for all remaining calculations. A more detailed analysis of the contribution of distant nuclear spins is presented in the following section.
2.6 Effect of Nuclear Spin Entanglement on
In this section, Model 4 is used to simulate the spin‐relaxation dynamics at G, which is the same strength used in the experiment[10.1038/s41467-023-44494-3]. Figs. 5(a) - (c) depict the variations in a population of polarization among a single electron spin, one of the three first-nearest nuclear spins, and the sum of two distant bath spins, respectively. These results indicate that the electron spin gradually transfers its polarization to all surrounding nuclei, including both nearest neighbors and the remote bath spins, as evidenced by the sharp increase in nuclear‐spin polarization in Figs. 5(b) and (c). As time progresses, the electron‐spin polarization continues to decay until full depolarization is achieved. Both the nearest-neighbor and the more distant nuclear spins exhibit an initial transient period (30 s) during which the nuclear spin polarization rises to positive values before decaying. After this transient regime, the entire bath enters a steady-state period, where similar dynamics are preserved over longer timescales, leading to a gradual buildup of polarization in the bath. At a later time, not achieved in our simulations, all spin populations will converge towards a value of 0.5, indicating that the spins are depolarizing and approaching thermal equilibrium. The fast coherent oscillation of the first nearest nitrogen nuclei, as exhibited in Fig. 5(d), can be explained by the oscillation between the and the states induced by the hyperfine interaction.
Now, turning to quantify the time, Fig. 5(a) showcases that the simulated polarization did not reach the thermal equilibrium point as the simulation has reached the allowed runtime in our supercomputer, which is 20 days. We anticipated that the simulation could take up to 80 days to reach the equilibrium point. To resolve this limitation, we simulated the polarization dynamics until it reached a maximum of 20 running days. Then we fitted the curve and extrapolated the rest of the behavior. This is still valid given that the polarization will, in principle, reach the equilibrium point at the polarization equal to 0.5. Furthermore, we have cross-checked with the polarization behavior at = 800 G as shown in Fig. 4(d) and found the equilibrium pattern. Note that the polarization can be simulated faster in the field close to GSLAC due to the narrower gap between and , as shown in Fig. 1. As a result, the fitted value of at = 90 G yields 694.77 s, which is quite consistent with the experimental value of 1 ms reported at 10 K in h10B15N[10.1038/s41467-023-44494-3]. For the small discrepancy, we hypothesize that this might arise from the absence of some correlations between different clusters, which are not captured even by our six-spin cluster approach. Moreover, while this work employed a ZFS value of 3.479 GHz, the ZFS increases to approximately 3.63 GHz at low temperatures [10.1038/s41467-021-24725-1]. The increased ZFS enhances the energy separation between the three states in the interval. Consequently, the time is expected to be slightly larger at low magnetic fields, bringing our results into closer agreement with the experiment.
We then systematically assess the impact of the evolution time on the fitted . As shown in Fig. 5(e), the converges once the simulation extends beyond 31.25 s, at which point the fitted value of 618.7 s underestimates the by 11% (compared with the at a maximum evolution time of 345.2 s, which requires 20 days of supercomputer runtime). To balance computational efficiency and accuracy in our subsequent investigation of as a function of magnetic field, we therefore limit all the remaining simulations to a 69 s evolution time, which likely incurs an artifact of approximately 4.51%.
2.7 Dependence on Magnetic Field
We now examine the variation of the time and its rate (1/) as a function of the field, as summarized in Figs. 6(a) and (b), respectively. Three distinct regimes can be identified: (i) the low‐field regime ( G), (ii) the near GSLAC regime ( G), and (iii) the high‐field regime ( G).
In low fields, decreases as approaches GSLAC near 1250 G. This can be attributed to the narrowing of the Zeeman splitting between and levels, as displayed in Fig. 1. The reduced energy gap enhances electron–nuclear flip–flop transitions mediated by the non‐secular hyperfine term , while the dipolar interaction term promotes faster nuclear–nuclear spin flip–flops by increasing bath fluctuations, as evidenced by the larger population changes of bath spins in the higher field shown in Supplementary S1. As a result, both mechanisms accelerate decoherence processes, leading to a shorter time with higher B strength. It is important to note that at lower fields ( G), the values presented in Fig. 6(a) are likely overestimated. This is because the current model reduces the Hilbert space by applying the reduced density matrix formalism to consider only the and sublevels. In practice, spin relaxation pathways involving flip-flop processes between and are also possible and should be taken into account. Accordingly, the total should be given by the relation
where and denote the relaxation times associated with spin flip-flop transitions between and , and between and , respectively. To address this limitation based on the same reduce density matrix, we computed the time under a field applied in the opposite direction. As shown in Fig. 6, the total is slightly lower than due to the presence of additional spin relaxation pathways. However, beyond a strength of 400 G, the total converges to values comparable to . This behavior arises because, at higher fields, the energy gap between and narrows while the gap between and widens, as illustrated in Fig. 1. Consequently, spin relaxation via the channel dominates, and the overall relaxation rate becomes effectively determined by this single pathway.
In the GSLAC region, the system exhibits different relaxation dynamics. Here, the and eigenstates undergo strong hybridization with nearby nuclear spin levels, resulting in oscillations in the population dynamics rather than the monotonic exponential decay, as shown in Supplementary S1. Under these conditions, a time cannot be extracted in the conventional sense based on an exponential fitting. This leads to an abrupt discontinuity in the 1/ behavior. This behavior has also been observed in 13C in NV center in diamond[10.1103/PhysRevB.101.155203], as well as 13C and 29Si in SiC[10.1038/s41524-021-00673-8].
In high fields, although nuclear–nuclear dipolar interactions persist and continue to contribute to bath-induced noise, their influence diminishes with increasing , as indicated by the reduced population fluctuations of bath spins in Supplementary S1. Furthermore, the increased energy separation between the and states suppresses hyperfine-mediated flip–flop transitions. Consequently, the overall relaxation rate decreases, leading to an enhancement of the time in this regime. It is important to note, however, that the computed values in this regime may still be slightly overestimated. This overestimation arises from the fact that, similar to the situation in low fields, the current model considers relaxation via the channel while neglecting potential relaxation through the pathway. As shown in Fig. 1, the Zeeman splitting between the and states increases substantially with . At sufficiently high fields, the energy gaps between and both states become approximately equal, making relaxation through both pathways equally probable. In this limit, the correct should be halved relative to the value obtained from a single-channel calculation; that is, .
For the field beyond 2000 G, this aspect remains a computational limitation in the present work, as indicated in Supplementary S1. This is due to the very large Zeeman splitting of and at high fields; thus, capturing the oscillation over such a wide energy range requires significantly increased computational resources. Addressing this in future studies would further refine the accuracy of predictions in the higher field regime.
In summary, while both nuclear–nuclear dipolar interactions and electron–nuclear hyperfine coupling contribute to spin relaxation across the entire range of magnetic fields, the dominant relaxation mechanism evolves as a function of the applied field strength. At low magnetic fields, both electron–nuclear and nuclear–nuclear interactions are enhanced, leading to faster relaxation and shorter times. In contrast, in high fields, increasing Zeeman splitting suppresses both types of flip–flop processes, resulting in weaker interactions and prolonged relaxation times. However, in the vicinity of the GSLAC, the system enters a strongly coupled regime where conventional descriptions based on exponential decay fail to capture the mechanism. This therefore results in a significantly reduced time.
3 Discussion
In this work, we have developed a spin dynamics model based on the cluster expansion method to investigate the relaxation mechanism of the V defect in bulk hBN. By systematically analyzing the roles of dipolar interactions and central spin–bath entanglement across four progressively refined models, we demonstrated that the V center represents a strongly coupled spin system with its three first-neighbor nitrogen nuclear spins. This necessitates the inclusion of all three nuclei within the central spin subsystem. As a result, the central spin model comprises four spins, while the spin bath is represented by additional distant, either moderately or weakly coupled 15N nuclear spins.
Our model was benchmarked against experimentally measured times, showing good agreement. The remaining discrepancies are likely due to the omission of inter-cluster correlations, which are not captured within the present 6-spin cluster expansion framework. Furthermore, we examined the dependence of on the external magnetic field and identified three distinct regimes: low-field, GSLAC, and high-field regions. The low- and high-field behaviors are well described by our model, while the dynamics near the GSLAC deviate from exponential relaxation due to strong electron–nuclear coupling, requiring more advanced treatments beyond the current framework.
While this work investigates the time in bulk hBN, dimensionality of the sample can also influence the results. In Model 4 of bulk hBN, the two distant nitrogen nuclei can sit either in the defected layer or in the neighboring layers. In contrast, in a monolayer hBN geometry, the two distant nitrogen nuclei would necessarily lie within the same plane and may therefore be located farther from the vacancy than in the bulk case. This would reduce their hyperfine interaction strength and, consequently, their contribution to decoherence. As a result, the nuclear spin bath-originated time in monolayer hBN is expected to be longer than in bulk or multilayer systems.
Overall, this work has established an effective and reliable model for describing relaxation dynamics in V centers in hBN. The insights gained into the underlying decoherence mechanisms provide a foundation for further exploration of spin dynamics in solid-state quantum systems. For instance, future studies could extend this framework to include additional relaxation mechanisms, such as interactions with nearby paramagnetic electron spins or boron nuclei, and investigate their contributions to the overall spin relaxation behavior. However, simply replacing the two distant 15N nuclei with 11B nuclei would increase the Hilbert space from 64 to 256, making such simulations computationally demanding with the current code. Improving computational efficiency will therefore be an essential step toward including boron nuclei in a systematic way. Additionally, the work may be continued by explicitly quantifying entanglement in these strongly coupled systems. A deep understanding of spin relaxation may lead to new strategies to engineer longer times, either through defect environment control or external manipulation. Such advancements would be valuable for optimizing V centers for applications in quantum sensing and quantum memory technologies.
4 Methods
4.1 Computational details
To solve the spin dynamics governed by the Lindblad master equation [Eq. (6)], we use our in-house MPI-parallelized C code implementing a fourth-order Runge-Kutta integration scheme. A timestep of 2.5 s is employed, as validated in Supplementary S4. To simulate the dynamics of V centers, we incorperate system-specific parameters, including the ZFS and HF interaction tensors. The ZFS is set to 3.479 GHz, as calculated using DFT together with the spin decotamination approach described in Ref.[10.1038/s41524-020-0305-x]. The HF interaction is obtained from real-space calculations, which include all nuclei within a 30-Å radius from the V defect in bulk hBN to resolve the finite size effect[10.1038/s42005-024-01668-9]. This corresponds to a total of 12,702 atoms in the supercell. The spin relaxation is then evaluated by monitoring the population of the V electron spin as it evolves toward thermal equilibrium. Here, the initial density matrix is prepared such that the V electron spin is polarized with 70% population in the state and 30% in the state, while all nuclear spins are initialized in the thermal equilibrium depolarized states.
Data availability
The data that support the findings of this study are available from the authors upon request.
Code Availability
The codes associated with this manuscript are available from the corresponding author upon request.
Acknowledgments
The work was supported by the National Research, Development and Innovation Office of Hungary (NKFIH) within the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004) and within the project FK 145395. This project is also funded by the European Commission within Horizon Europe projects (Grant Nos. 101156088 and 101129663). This research is part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy- EXC-2111-390814868 (MCQST) and as part of the CRC 1375 NOA project C2. The authors acknowledge support from the Federal Ministry of Research, Technology and Space (BMFTR) under grant number 13N16292 (ATOMIQS). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at Leibniz Supercomputing Centre (www.lrz.de) and on its Linux-Cluster.
Author contributions
C.C. initiated the study and carried out the simulations. The manuscript was written by C.C. with inputs from T.V. and V.I. The work was supervised with T.V. and V.I.
Competing interests
The authors declare no competing interests.